1.In the circuit below, if VCC = 5V, ?? = 4?Ω, ?1 = 60?Ω, ?2 = 40?Ω, ?? = 3?Ω, ?? = 4?Ω, ? = 75, ?? = 80?, ??? (??) = 0.7?

a. What kind of amp is the circuit? Why is that? What are its features?
b. Find the VECQ and ICQ values. Draw the load line.
c. Draw the small signal curve circuit of the circuit. Find all the intermediate steps of your analysis and find the voltage gain, input and output resistors.

Į REŽ R
Answer #1

(a) common collector amplifier used between amplifier stages for impedance matching. It has high current gain and low Voltage gain.

DC Analysis: • open circuit all capacitors and draw the ciocult A p5V 60k Zakr -0.7 +YI cokr HD Thevenin equivalent between Ae ca akn yte = (+1) IG -3V of VCEO 2012 m Apply KUE COEB lopp. 418 + 3KR XIO (p+1) + 0.7 +26+R(IB) - 43 V = 0 IB = 2.30 3kx76Transconductance 980 = TE vp = Theron al voltage : = gomv at room temperature gos. 3 0.685mA 26mu som = 26-68 ms. ] load line(c) small segonal Analysies.. T5 - = 2.81 km 26-66 ms Reb v At out put sede Yo = VA 800 =lis. 2 KR 0.693 mA = rollrellau) xf@2.52 kr Vo = — xus 2052kr +482 V = 0.387 Vs - 0.383 VS 208kr substitute ② 00387 xks Vo = -128.3 76KXS - 2.81kr Voltage galos